A bulb manufacturer claims that the lives of its bulbs are normally distributed with a mean of 6000 hours and a standard deviation of 400 hours. A random sample of 16 bulbs had an average life of 5850 hours. If the manufacturer’s claim is correct-

a. What is the sampling distribution of the sample mean?

b. what is the probability of finding a sample mean of 5850 or less?


1
Expert's answer
2022-05-12T04:08:28-0400

a. If the population is normally distributed, then the sampling distribution of "\bar{X}" is normally distributed for any sample size "n."

Let "\bar{X}=" the sample mean: "\bar{X}\sim N(\mu, \sigma^2/n)."

Given "\mu=600h, \sigma=400h, n=16"


"\bar{X}\sim N(6000, 400^2/16)"

b.


"P(\bar{X}\le5850)=P(Z\le\dfrac{5850-6000}{400/\sqrt{16}})"

"=P(Z\le-1.5)\approx0.0668"